# Data-Efficient Extremum-Seeking Control Using Kernel-Based Function Approximation

### v1

@@ 3 Data-efficient extremum-seeking using surrogate modeling
This section consists of three parts. In Section 3.1, we outline how an approximation of the steady-state input-output map can be used to perform a parameter update step (instead of a standard update step as in Algorithm 1) whenever a decrease in cost can be guaranteed, to reduce the number of experiments needed to optimize performance. In Section 3.2, we discuss how such approximation can be obtained from data collected during standard parameter update steps. Finally, in Section 3.3 we discuss how the bounds used to assess whether a decrease can be guaranteed are obtained and present the full approach.
@@ 3.1 Parameter updates based on cost function approximation
Suppose that a continuously differentiable approximation $m(\cdot)$ of the steady-state input-output map $f(\cdot)$ described in Section 2 is available. In case $-\nabla m(\theta)$ is a descent direction of $f(\theta)$, i.e., if
@@
an update of the form
$\theta^{+}=\theta-\mu\nabla m(\theta)$ (20)
guarantees a decrease in $f(\theta)$ provided that the optimizer gain $\mu\in\mathbb{R}_{>0}$ is suitably chosen. However, unlike the standard parameter update (17) in Algorithm 1, an update of the form (20) does not require additional experiments to be performed to determine the search direction. Hence, performing updates of the form (20) reduces the total number of experiments that are needed to solve the optimization problem in (7), thus saving time.
A well-known condition used in static optimization to determine if an optimizer gain $\mu$ guarantees a decrease in $f(\theta)$ during a parameter update step of the form (20) is the so-called Armijo condition (see, e.g., Nocedal and Wright (2006, Chapter 3)), given by
where $\theta^{+}$ is as in (20) and $c\in(0,1)$ is a control parameter. If $-\nabla m(\theta)$ is a descent direction (cf. (19)), satisfaction of the Armijo condition (21) guarantees a decrease in $f(\theta)$ as a result of the update step, and convergence to the optimizer if it is satisfied for all update steps (provided that Assumption 2(iii) holds, cf. Remark 3). Hence, a suitable optimizer gain $\mu$ for which a decrease in $f(\theta)$ is guaranteed can for example be obtained by performing a backtracking line search until a $\mu$ is found for which (21) is satisfied.
However, in contrast to static optimization where the cost function (and its gradient) are typically known, the steady-state input-output map $f(\cdot)$ is unknown in the context of ESC, and can only be evaluated approximately by performing experiments. Moreover, its gradient $\nabla f(\cdot)$ is unavailable, and can only be approximated by performing multiple (approximate) evaluations of $f(\cdot)$. Hence, verifying satisfaction of (19) and (21) (to determine if $-\nabla m(\theta)$ is a descent direction and to find a suitable $\mu$) would require performing multiple experiments, potentially adversely affecting the benefit of performing update steps of the form (20) in the first place. Yet, suppose two bounds $\underline{b}(\theta)$ and $\overline{b}(\theta,\mu)$ are available that can be evaluated without performing experiments and that satisfy
$\nabla f(\theta)^{\intercal}\nabla m(\theta)\geq\underline{b}(\theta)$ (22)
@@
$f(\theta^{+})-f(\theta)+c\mu\nabla f(\theta)^{\intercal}\nabla m(\theta)\leq\overline{b}(\theta,\mu)$ (23)
with $\theta^{+}$ as in (20). Then, guaranteed satisfaction of (19) and (21) can be verified by checking whether $\underline{b}(\theta)>0$ and $\overline{b}(\theta,\mu)\leq 0$.
Given the discussion above, Figure 1 provides an illustration of an approach aimed at reducing the total number of experiments needed to solve the optimization problem in (7) in the context of ESC, which we briefly describe next. More details will be provided in Section 3.3.
Suppose that a dataset $\mathcal{D}_{k}=\{(\widetilde{\theta}_{i},\widetilde{y}_{i}):i=1,\dots,N_{k}\}$ of $N_{k}$ inputs $\widetilde{\theta}_{i}$ and corresponding measured outputs $\widetilde{y}_{i}$, e.g., as in (14) and (15), respectively, is available. Such dataset can for example be obtained by performing a single standard parameter update step as in Algorithm 1. From $\mathcal{D}_{k}$, we construct an approximation $m_{k}(\cdot)$ of the steady-state input-output map $f(\cdot)$ using kernel-based function approximation, as we will explain in Section 3.2. For such kernel-based approximation, it is also possible to derive bounds $\underline{b}(\theta)$ and $\overline{b}(\theta,\mu)$ satisfying (22) and (23) that can be evaluated without performing additional experiments, as we will show in Section 3.3. Hence, it can be verified without additional experiments whether $-\nabla m_{k}(\theta)$ is a descent direction of $f(\theta)$ by checking whether $\underline{b}(\theta)>0$ (cf. (19) and (22)). If this is the case, we perform a backtracking line search over $\mu$ to obtain a suitable optimizer gain $\mu$ for which $\overline{b}(\theta,\mu)\leq 0$, such that the Armijo condition (21) is guaranteed to be satisfied (cf. (23)). If such $\mu$ is found, we perform the parameter update step as in (20). In case $\underline{b}(\theta)\leq 0$ or no $\mu$ satisfying $\overline{b}(\theta,\mu)\leq 0$ has been found during the line search, we instead perform a standard parameter update step as in Algorithm 1 to: (i) update the parameters $\theta$; and (ii) collect more data that are added to the dataset $\mathcal{D}_{k+1}$ to construct a new approximation $m_{k+1}(\cdot)$ during the next iteration of the ESC algorithm.
@@
We will refer to the approach outlined above as kernel-based extremum-seeking control (KB-ESC). Here, ‘kernel-based’ refers to the fact that a kernel-based approximation $m_{k}(\cdot)$ of the input-output map $f(\cdot)$ is constructed and used for some parameter update steps. Furthermore, we will refer to parameter update steps as in (20) as kernel-based update steps, to distinguish them from the standard parameter update steps described in Algorithm 1.
@@ 3.2 Approximating the steady-state input-output map using kernels
Given a collected dataset $\mathcal{D}_{k}=\{(\widetilde{\theta}_{i},\widetilde{y}_{i}):i=1,\dots,N_{k}\}$ consisting of $N_{k}$ constant inputs $\widetilde{\theta}_{i}$ and corresponding measured outputs $\widetilde{y}_{i}$, e.g., as in (14) and (15), respectively, we choose to construct the approximation $m_{k}(\cdot)$ of the input-output map $f(\cdot)$ using kernel-based function approximation. We take a kernel to be a continuous, symmetric function $\kappa:\mathbb{R}^{n_{\theta}}\times\mathbb{R}^{n_{\theta}}\to\mathbb{R}$ that is positive definite according to the following definition (see, e.g., Schölkopf et al. (2001, Definition 3)).
@@
By the Moore-Aronszajn theorem (Aronszajn, 1950), any kernel satisfying Definition 6 has a unique reproducing kernel Hilbert space (RKHS) of functions associated with it. This RKHS, which we will denote by $\mathcal{H}$, is the completion of the space of functions of the form $g(\cdot)=\sum_{i=1}^{N}$, $\alpha_{i}\kappa(\cdot,\theta_{i})$ with respect to the norm $\lVert g\rVert:=\sqrt{\langle g,g\rangle}$ corresponding to the inner product $\langle g_{1},g_{2}\rangle=\left\langle\sum_{i=1}^{N}\alpha_{i}\kappa(\cdot,\theta_{i}),\sum_{j=1}^{M}\beta_{j}\kappa(\cdot,\theta_{j}^{\prime})\right\rangle=\sum_{i=1}^{N}\sum_{j=1}^{M}\alpha_{i}\beta_{j}\kappa(\theta_{i},\theta_{j}^{\prime})$. Here, $N,M\in\mathbb{N}$, $\alpha_{1},\dots,\alpha_{N},\allowbreak\beta_{1},\dots,\beta_{M}\in\mathbb{R}$, and $\theta_{1},\dots,\theta_{N},\theta^{\prime}_{1},\dots,\theta^{\prime}_{M}\in\mathbb{R}^{n_{\theta}}$. Note that $\lVert g\rVert^{2}$ takes the quadratic form
@@
In order to construct an approximation $m_{k}(\cdot)$ of the steady-state input-output map $f(\cdot)$ on the basis of the dataset $\mathcal{D}_{k}$, and draw conclusions on the accuracy of the approximation, we adopt the following two assumptions whose motivation will become clear shortly (see also Scharnhorst et al. (2023, Assumptions 1 and 2)).
@@
An upper bound $\bar{\delta}\in\mathbb{R}_{\geq 0}$ such that $\lvert f(\widetilde{\theta}_{i})-\widetilde{y}_{i}\rvert\leq\bar{\delta}$ for all measured outputs $\widetilde{y}_{i}$, $i=1,\dots,N_{k}$, is known, i.e., for every measured output $\widetilde{y}_{i}$ its distance to the corresponding steady-state value $f(\widetilde{\theta}_{i})$ is at most $\bar{\delta}$.
@@
Note that given Assumptions 2(iv) and 2(vi), Assumption 7 can be satisfied for any $\bar{\delta}\in\mathbb{R}_{>0}$ by choosing a sufficiently long waiting time $T$. Furthermore, the choice of kernel $\kappa$ in Assumption 8 allows including prior knowledge that might be available about the input-output map $f(\cdot)$. In case no prior knowledge about this input-output map is available, universal kernels such as the squared exponential kernel
@@
with $\sigma$ a tunable length scale, could be chosen since such kernels have the ability to approximate any continuous function arbitrarily closely on a compact set (see, e.g., Micchelli et al. (2006)).
Given Assumptions 7 and 8, we construct the approximation $m_{k}(\cdot)$ on the basis of the dataset $\mathcal{D}_{k}$ by solving the optimization problem
$m_{k}(\cdot)=\underset{m\in\mathcal{H}}{\rm argmin}$ $\>\lVert m\rVert^{2}$ (26a)
s.t. $\>\lvert m(\widetilde{\theta}_{i})-\widetilde{y}_{i}\rvert\leq\bar{\delta}\quad\forall i\in\{1,\dots,N_{k}\},$ (26b)
@@
Note that (26) is guaranteed to have a solution because the feasible set is non-empty, as by Assumptions 7 and 8 it contains the steady-state input-output map $f(\cdot)$. Moreover, note that (26) is equivalent to
$m_{k}(\cdot)=\underset{m\in\mathcal{H}}{\rm argmin}\>\left(\lVert m\rVert^{2}+\sum_{i=1}^{N_{k}}\zeta_{\bar{\delta}}(m(\widetilde{\theta_{i}})-\widetilde{y}_{i})\right)$ (27)
with
$\zeta_{\bar{\delta}}(z):=\begin{cases}0,&\text{if }\lvert z\rvert\leq\bar{\delta},\\ \infty,&\text{otherwise},\end{cases}$ (28)
which allows application of the representer theorem (Schölkopf et al., 2001, Theorem 1). The representer theorem states that the solution to (26) has the form
@@
where $K_{\Omega}(\cdot):=[\kappa(\cdot,\widetilde{\theta}_{1})\,\cdots\,\kappa(\cdot,\widetilde{\theta}_{N_{k}})]$ is a row vector of kernel functions centered at the inputs $\Omega=\{\widetilde{\theta}_{1},\dots,\widetilde{\theta}_{N_{k}}\}$. Hence, substituting a solution $m(\cdot)=K_{\Omega}(\cdot)\alpha$ in (26), and using (24), we can reformulate the optimization problem as the quadratic programming problem
$\alpha=\underset{\bar{\alpha}\in\mathbb{R}^{N_{k}}}{\rm argmin}\>$ $\bar{\alpha}^{\intercal}K_{\Omega\Omega}\bar{\alpha}$ (30a)
s.t. $\lvert K_{\Omega}(\widetilde{\theta}_{i})\bar{\alpha}-\widetilde{y}_{i}\rvert\leq\bar{\delta}\quad\forall i\in\{1,\dots,N_{k}\}.$ (30b)
As a consequence, every time a new dataset $\mathcal{D}_{k}$ is obtained by adding new data pairs $(\widetilde{\theta},\widetilde{y})$ to the previous dataset, a new approximation $m_{k}(\cdot)$ can be obtained by simply solving (30) to obtain a new weight vector $\alpha\in\mathbb{R}^{N_{k}}$ and using (29). Furthermore, since the weight vector $\alpha$ only depends on the dataset $\mathcal{D}_{k}$ and the chosen kernel $\kappa$, the gradient of the approximation $m_{k}(\cdot)$ is readily available as (cf. (29))
@@
denotes a matrix of partial derivatives of the kernel functions centered at the inputs $\Omega=\{\widetilde{\theta}_{1},\dots,\widetilde{\theta}_{N_{k}}\}$.

### v2

@@ 3 Data-efficient extremum-seeking using function approximation
This section consists of three parts. In Section 3.1, we outline how an approximation of the steady-state input-output map can be used to perform a parameter update step (instead of a standard update step as in (10)-(13)) whenever a decrease in cost can be guaranteed, to reduce the number of measurements needed to optimize performance. In Section 3.2, we discuss how such an approximation can be obtained from data collected during regular operation of the extremum-seeking controller. Finally, in Section 3.3 we discuss how the bounds used to assess whether a decrease can be guaranteed are obtained and present the full approach.
@@ 3.1 Parameter updates based on cost function approximation
Suppose that a continuously differentiable approximation $m(\cdot)$ of the steady-state input-output map $f(\cdot)$ described in Section 2 is available. We will describe in Section 3.2 how such an approximation can be obtained on the basis of input-output data collected during regular operation of the extremum-seeking controller. Since $m(\cdot)$ approximates $f(\cdot)$, its gradient $\nabla m(\theta)$ also forms an approximation of the gradient of $f(\theta)$. Hence, we can use it for a gradient-based update of $\theta$ aimed at minimizing $f(\theta)$. That is, we can use $\nabla m(\theta)$ for an update of the form
$\theta^{+}=\theta-\mu\nabla m(\theta)$ (19)
with $\mu\in\mathbb{R}_{>0}$ a suitably chosen optimizer gain. Unlike the standard parameter update (10)-(13), an update of the form (19) does not require additional measurements to be performed. Hence, replacing standard updates by updates of the form (19) reduces the total number of (time-consuming and costly) measurements required to solve the optimization problem in (7).
A common approach for choosing the optimizer gain $\mu$ in (19), used in static optimization, is to perform a backtracking line search until the so-called Armijo condition is satisfied (see, e.g., Nocedal and Wright (2006, Chapter 3)). This condition is given by
where $\theta^{+}$ is as in (19) and $c\in(0,1)$ is a standard parameter that is typically chosen in the order of $10^{-4}$ (see, e.g., Nocedal and Wright (2006, Chapter 3)). If $-\nabla m(\theta)$ is a descent direction of $f(\theta)$, i.e., if
@@
then satisfaction of the Armijo condition (20) guarantees a decrease in $f(\theta)$ as a result of the update step. Furthermore, if (20) and (21) are satisfied for all update steps, $\theta$ converges to the optimizer (provided that Assumption 2(iii) holds, cf. Remark 3).
However, in contrast to static optimization, where the cost function (and its gradient) are typically known, the steady-state input-output map $f(\cdot)$ is unknown in the context of ESC, and can only be evaluated approximately by performing measurements. Moreover, its gradient $\nabla f(\cdot)$ is unavailable, and can only be approximated by performing multiple (approximate) evaluations of $f(\cdot)$. Hence, verifying satisfaction of (20) and (21) (to determine if $-\nabla m(\theta)$ is a descent direction and to find a suitable $\mu$) would require performing multiple measurements, potentially adversely affecting the benefit of performing update steps of the form (19) in the first place. Yet, we will show in Section 3.3 that when the approximation $m(\cdot)$ is constructed in the way that we will describe in Section 3.2, it is possible to derive bounds $\underaccent{\bar}{b}(\theta)$ and $\bar{b}(\theta,\mu)$ that can be evaluated without performing measurements and that satisfy
$\nabla f(\theta)^{\intercal}\nabla m(\theta)\geq\underaccent{\bar}{b}(\theta)$ (22)
@@
$f(\theta^{+})-f(\theta)+c\mu\nabla f(\theta)^{\intercal}\nabla m(\theta)\leq\bar{b}(\theta,\mu)$ (23)
with $\theta^{+}$ as in (19). Hence, we can use these bounds to verify whether (20) and (21) are satisfied, without performing measurements, by checking whether $\underaccent{\bar}{b}(\theta)>0$ and $\bar{b}(\theta,\mu)\leq 0$.
Given the discussion above, we provide in Figure 1 an illustration of our approach aimed at reducing the total number of measurements needed to solve the optimization problem in (7) in the context of ESC, which we briefly describe next. More details will be provided in Section 3.3.
@@
Suppose that a dataset $\mathcal{D}_{k}=\{(\widetilde{\theta}_{i},\widetilde{y}_{i})\mid i=1,2,\dots,N_{k}\}$ of $N_{k}$ inputs $\widetilde{\theta}_{i}$ and corresponding measured outputs $\widetilde{y}_{i}$, e.g., as in (10) and (11), respectively, is available. If such a dataset is not available a priori, we simply obtain it during a single standard parameter update step as in (10)-(13). From $\mathcal{D}_{k}$, we construct an approximation $m_{k}(\cdot)$ of the steady-state input-output map $f(\cdot)$ using kernel-based function approximation, as we will explain in Section 3.2. Then, for the current input $\theta$, we verify whether $-\nabla m_{k}(\theta)$ is a descent direction of $f(\theta)$ by checking whether $\underaccent{\bar}{b}(\theta)>0$ (cf. (21) and (22)). If this is the case, we perform a backtracking line search over $\mu$ to obtain a suitable optimizer gain $\mu$ for which $\bar{b}(\theta,\mu)\leq 0$, such that the Armijo condition (20) is guaranteed to be satisfied (cf. (23)). If such a $\mu$ is found, we perform the parameter update step as in (19), which does not require additional measurements. In case $\underaccent{\bar}{b}(\theta)\leq 0$ or no $\mu$ satisfying $\bar{b}(\theta,\mu)\leq 0$ has been found during the line search, we instead perform a standard parameter update step as in (10)-(13) to update the parameters $\theta$. Subsequently, we add the input-output data collected during this standard update step to the dataset $\mathcal{D}_{k+1}$ to construct an updated approximation $m_{k+1}(\cdot)$ during the next iteration of the ESC algorithm.
We will refer to the approach outlined above as kernel-based extremum-seeking control (KB-ESC). Here, ‘kernel-based’ refers to the fact that a kernel-based approximation $m_{k}(\cdot)$ of the input-output map $f(\cdot)$ is constructed and used for some parameter update steps. Furthermore, we will refer to parameter update steps as in (19) as kernel-based update steps, to distinguish them from the standard parameter update steps in (10)-(13).
@@ 3.2 Approximating the steady-state input-output map using kernels
As mentioned in Section 3.1, we choose to construct the approximation $m_{k}(\cdot)$ of the input-output map $f(\cdot)$ using kernel-based function approximation. We take a kernel to be a continuous, symmetric function $\kappa:\mathbb{R}^{n_{\theta}}\times\mathbb{R}^{n_{\theta}}\to\mathbb{R}$ that is positive definite according to the following definition (see, e.g., Schölkopf et al. (2001, Definition 3)).
@@
By the Moore-Aronszajn theorem (Aronszajn, 1950), any kernel satisfying Definition 6 has a unique reproducing kernel Hilbert space (RKHS) of functions associated with it. This RKHS, which we will denote by $\mathcal{H}$, is the completion of the space of functions of the form $g(\cdot)=\sum_{i=1}^{N}\alpha_{i}\kappa(\cdot,\theta_{i})$ with respect to the norm $\lVert g\rVert:=\sqrt{\langle g,g\rangle}$ corresponding to the inner product $\langle g_{1},g_{2}\rangle=\left\langle\sum_{i=1}^{N}\alpha_{i}\kappa(\cdot,\theta_{i}),\sum_{j=1}^{M}\beta_{j}\kappa(\cdot,\theta_{j}^{\prime})\right\rangle=\sum_{i=1}^{N}\sum_{j=1}^{M}\alpha_{i}\beta_{j}\kappa(\theta_{i},\theta_{j}^{\prime})$. Here, $N,M\in\mathbb{N}$, $\alpha_{1},\dots,\alpha_{N},\allowbreak\beta_{1},\dots,\beta_{M}\in\mathbb{R}$, and $\theta_{1},\dots,\theta_{N},\theta^{\prime}_{1},\dots,\theta^{\prime}_{M}\in\mathbb{R}^{n_{\theta}}$. Note that $\lVert g\rVert^{2}$ takes the quadratic form
@@
We adopt the following two assumptions (see also Scharnhorst et al. (2023, Assumptions 1 and 2)).
@@
An upper bound $\bar{\delta}\in\mathbb{R}_{\geq 0}$ is known, such that $\lvert f(\widetilde{\theta}_{i})-\widetilde{y}_{i}\rvert\leq\bar{\delta}$ for all measured outputs $\widetilde{y}_{i}$, $i=1,2,\dots,N_{k}$. That is, for every measured output $\widetilde{y}_{i}$, its distance to the corresponding steady-state value $f(\widetilde{\theta}_{i})$ is at most $\bar{\delta}$.
@@
These assumptions will allow us to derive the bounds $\underaccent{\bar}{b}(\theta)$ and $\bar{b}(\theta,\mu)$ mentioned in Section 3.1, which we will use to verify, without performing measurements, whether $-\nabla m_{k}(\theta)$ is a descent direction of $f(\theta)$ and whether the Armijo condition (20) is satisfied for a given $\mu$. That is, whether we can perform a kernel-based update step with this $\mu$.
@@
Note that, instead of requiring $\bar{\delta}$ to be known in Assumption 7, it can also be considered a design parameter since, given Assumptions 2(iv) and 2(vi), Assumption 7 can be satisfied for any $\bar{\delta}\in\mathbb{R}_{>0}$ by choosing a sufficiently long waiting time $T$. Furthermore, the choice of kernel $\kappa$ in Assumption 8 allows including prior knowledge that might be available about the input-output map $f(\cdot)$. In case no prior knowledge about this input-output map is available, universal kernels such as the squared exponential kernel
@@
with $\sigma$ a tunable length scale, could be chosen since such kernels have the ability to approximate any continuous function arbitrarily closely on a compact set (see, e.g., Micchelli et al. (2006)). The bound $\Gamma$ in Assumption 8 can be seen as a design parameter, which should be chosen large enough to guarantee that $\Gamma\geq\lVert f\rVert$, similar to how in standard sampled-data ESC the waiting time $T$ is a design parameter that should be chosen large enough to guarantee the system output $y$ is close to being in steady state when the output is measured.
Given Assumptions 7 and 8, we construct the approximation $m_{k}(\cdot)$ on the basis of data $\mathcal{D}_{k}=\{(\widetilde{\theta}_{i},\widetilde{y}_{i})\mid i=1,2,\dots,N_{k}\}$ collected, e.g., during standard update steps (10)-(13), by solving the optimization problem
$m_{k}(\cdot)=\operatorname*{\arg\min}_{m\in\mathcal{H}}$ $\>\lVert m\rVert^{2}$ (26a)
s.t. $\>\lvert m(\widetilde{\theta}_{i})-\widetilde{y}_{i}\rvert\leq\bar{\delta}$
$\quad\forall i\in\{1,2,\dots,N_{k}\},$ (26b)
@@
Note that, by Assumptions 7 and 8, the steady-state input-output map $f(\cdot)$ is a feasible solution to (26), and thus (26) is guaranteed to have a solution.
By the representer theorem (see, e.g., Schölkopf et al. (2001, Theorem 1)) the solution to (26) has the form
@@
where $K_{\Omega}(\cdot):=[\kappa(\cdot,\widetilde{\theta}_{1})\,\kappa(\cdot,\widetilde{\theta}_{2})\,\cdots\,\kappa(\cdot,\widetilde{\theta}_{N_{k}})]$ is a row vector of kernel functions centered at the inputs $\Omega=\{\widetilde{\theta}_{1},\widetilde{\theta}_{2},\dots,\widetilde{\theta}_{N_{k}}\}$. Hence, substituting a solution $m(\cdot)=K_{\Omega}(\cdot)\alpha$ in (26), and using (24), shows that every time a new dataset $\mathcal{D}_{k}$ is obtained by adding new data pairs $(\widetilde{\theta},\widetilde{y})$ to the previous dataset, a new approximation $m_{k}(\cdot)$ can be constructed by obtaining a new weight vector $\alpha\in\mathbb{R}^{N_{k}}$ by simply solving the quadratic programming problem
$\alpha=\operatorname*{\arg\min}_{\bar{\alpha}\in\mathbb{R}^{N_{k}}}\>$ $\bar{\alpha}^{\intercal}K_{\Omega\Omega}\bar{\alpha}$ (28a)
s.t. $\lvert K_{\Omega}(\widetilde{\theta}_{i})\bar{\alpha}-\widetilde{y}_{i}\rvert\leq\bar{\delta}$
$\quad\forall i\in\{1,\dots,N_{k}\}$ (28b)
and using (27). Furthermore, since the weight vector $\alpha$ only depends on the dataset $\mathcal{D}_{k}$ and the chosen kernel $\kappa$, the gradient of the approximation $m_{k}(\cdot)$ is readily available as (cf. (27))
@@
denotes a matrix of partial derivatives of the kernel functions centered at the inputs $\Omega=\{\widetilde{\theta}_{1},\widetilde{\theta}_{2},\dots,\widetilde{\theta}_{N_{k}}\}$.